# On the Design and Analysis of LLM-Based Algorithms

### v1

@@ Abstract
We initiate a formal investigation into the design and analysis of LLM-based algorithms, i.e. algorithms that contain one or multiple calls of large language models (LLMs) as sub-routines and critically rely on the capabilities of LLMs. While LLM-based algorithms, ranging from basic LLM calls with prompt engineering to complicated LLM-powered agent systems and compound AI systems, have achieved remarkable empirical success, the design and optimization of them have mostly relied on heuristics and trial-and-errors, which is largely due to a lack of formal and analytical study for these algorithms. To fill this gap, we start by identifying the computational-graph representation of LLM-based algorithms, the design principle of task decomposition, and some key abstractions, which then facilitate our formal analysis for the accuracy and efficiency of LLM-based algorithms, despite the black-box nature of LLMs. We further consider parallel decomposition for a case study, providing extensive analytical and empirical study for four concrete examples of this pattern. Our proposed framework holds promise for advancing LLM-based algorithms, by revealing the reasons behind curious empirical phenomena, guiding the choices of hyperparameters, predicting the empirical performance of algorithms, and inspiring new algorithm design. To promote further study of LLM-based algorithms, we release our source code at https://github.com/modelscope/agentscope/tree/main/examples/paper_llm_based_algorithm.
@@ 1 Introduction
The rapid developments of pre-trained large language models (LLMs) in the past few years [BCE+23, WTB+22, SMK23] have given rise to the new paradigm of algorithmic problem solving by utilizing LLMs as general-purpose solvers and prompting them with task descriptions plus additional inputs that can help enhance their performance [WWS+22, KGR+22, ZYY24]. Meanwhile, it is also well recognized that even the best LLMs today still exhibit various limitations, such as finite context window sizes and difficulty with complex reasoning. Some of these limitations are fundamental [MS23, PNP24, TTC+24, HR24], and resolving them requires new breakthroughs. Moreover, in resource-constrained scenarios where using the state-of-the-art LLMs is not feasible, one might have to resort to smaller and weaker LLMs for solving complex tasks.
All these have motivated the developments of what we call LLM-based algorithms, namely, algorithms that contain one or multiple LLM calls as sub-routines and fundamentally rely on the capabilities of LLMs. In its most basic form, an LLM-based algorithm can be a combination of one or multiple LLM calls and some prompt engineering [YYZ+23, BBK+23, LLLD23, SLC+23, PKH+24, CLGI24]. More advanced examples include LLM-powered agent systems [MDL+23, WBZ+23, HZC+23, GLP+24, SST+23, ZWK+24, PKB+24, XCG+23] and compound AI systems [ZKC+24, CDH+24] that augment LLMs with additional abilities like tool use and long-term memory, as well as the emerging paradigm of LLM programming [SSC+23, KTF+23, KSM+24, ZYX+24].
@@
  • As a case study, in Section 3 we provide an in-depth study of parallel decomposition, a basic pattern of task decomposition and a building block for more sophisticated LLM-based algorithms. After introducing the general formulation and analysis for this pattern, we present our algorithm design and analysis for four diverse tasks, which are further validated by extensive numerical experiments.
@@ 2 LLM-based algorithms
  • Other classes of problems might have limited problem sizes, but require complex reasoning for solving them correctly. In this case, LLM-based algorithms can benefit substantially from sequential task decomposition such as step-by-step thinking and acting [YZY+23], or parallel decomposition like generating multiple reasoning paths and then aggregating them for a better solution [SLC+23].
@@ 2.1 Algorithms as computational graphs
Given such nodes, the computational graph is built by connecting them with directed edges that specify the data flows within the LLM-based algorithm. For example, Figure 2(a) shows the graph for the pattern of parallel decomposition, which divides the input problem into parallel sub-tasks, solves each of them with one LLM node, and aggregates the results for the final solution; Figure 2(b) illustrates an algorithm for book-length summarization [CLGI24, Ope24b], which divides the input text into multiple chunks and maintains a global summary via incremental updating; and Figure 2(c) represents the ReAct algorithm [YZY+23], which consists of multiple iterations of reasoning, tool use and aggregation. The computational-graph formulation is expressive enough to cover these diverse LLM-based algorithms that have been proposed separately in prior works, and facilitates formal analysis for all of them in a unified manner. Configurations of an LLM-based algorithm include how task decomposition is done (reflected by the graph topology and the sub-tasks of graph nodes), the methods of prompting and parsing for each LLM node, configurations of the LLM(s) being used, and so on. The LLM nodes within one graph might use the same or different backbone LLMs.
@@ 2.2 Key abstractions
Error metrics are task-specific in general. Moreover, the overall algorithm and the LLM nodes within it might have different error metrics. For instance, consider an LLM-based algorithm that answers a given question by first retrieving relevant sentences from the input text with multiple LLM nodes, and then reasoning over the extracted sentences with another LLM node. In this case, the overall algorithm is evaluated by whether it answers the question correctly, while each LLM node responsible for retrieval is evaluated by whether it successfully extracts all necessary sentences within the input text that it handles.
@@
Assumptions on LLM inference service, on the other hand, are task-agnostic and only affect the cost metrics, not error metrics. In particular, they determine the relation between the cost metrics and the lengths of prompt and generated text for each LLM call, the parallelism of multiple LLM calls, among others. While LLM inference service in practice can be very diverse, we will see in the next subsection that unified and formal analysis is possible with appropriate abstractions of it.
@@ 2.3 Formal analysis: accuracy and efficiency
$\mathcal{C}$ $\leq\mathcal{C}(\text{prefilling})+\mathcal{C}(\text{decoding})=\mathcal{C}_{\textsf{pre}}(\mathcal{L}_{\textsf{pre}})+\mathcal{C}_{\textsf{dec}}(\mathcal{L}_{\textsf{pre}},\mathcal{L}_{\textsf{dec}}).$ (2)
@@
  • Latency: if LLM inference is memory-bound, then $\mathcal{C}_{\textsf{pre}}(\mathcal{L}_{\textsf{pre}})=O(1)$ and $\mathcal{C}_{\textsf{dec}}(\mathcal{L}_{\textsf{pre}},\mathcal{L}_{\textsf{dec}})=O(\mathcal{L}_{\textsf{dec}})$; on the other hand, in the compute-bound scenario, $\mathcal{C}_{\textsf{pre}}$ and $\mathcal{C}_{\textsf{dec}}$ might grow linearly or even quadratically (e.g. for Transformers with full attention [VSP+17]) with $\mathcal{L}_{\textsf{pre}}$ and $\mathcal{L}_{\textsf{dec}}$.
@@ 3 Case study: parallel decomposition
This section focuses on parallel decomposition, a basic pattern visualized in Figure 2(a). An algorithm of this pattern first divides the input problem into multiple independent sub-tasks, solves each of them with one LLM node, and finally aggregates the results with an LLM or non-LLM node for the final solution. The intermediate sub-tasks can be solved sequentially or in parallel, which has impacts on certain cost metrics of the overall algorithm. This basic form of task decomposition can be used as a building block for more sophisticated algorithms. Despite its simplicity, interesting analysis and a wide variety of concrete tasks and algorithms can be derived from this pattern, which are presented in the remaining of this section.
@@ 3.1 Notations and analysis
  • Finally, let $\mathcal{L}_{\textsf{sys}}$ be an upper bound for the length of the system prompt of each LLM call. The presence of $\mathcal{L}_{\textsf{sys}}$, which can be large in practice, is essentially due to the fact that the LLM is used as a general-purpose sub-routine, hence specifying the concrete task for each LLM call constitutes part of the complexity.
@@
For notational convenience, we will often write $f(n)\lesssim g(n)$ in place of $f(n)=O(g(n))$, which means there exists a universal constant $C>0$ such that $f(n)\leq C\cdot g(n)$ for any positive integer $n$. In addition, $f(n)\asymp g(n)$ means $f(n)\lesssim g(n)$ and $g(n)\lesssim f(n)$ both hold. To further simplify notation, we will often omit the big-O notation in the input variables of cost functions $\mathcal{C}_{\textsf{pre}}$ and $\mathcal{C}_{\textsf{dec}}$; for example, given a prompt of length $\mathcal{L}_{\textsf{sys}}+O(n)$, we will write the cost of the prefilling phase as $\mathcal{C}_{\textsf{pre}}(\mathcal{L}_{\textsf{sys}}+n)$ rather than $\mathcal{C}_{\textsf{pre}}(\mathcal{L}_{\textsf{sys}}+O(n))$.
@@
$\leq k\times\Big{(}\mathcal{C}_{\textsf{pre}}(\mathcal{L}_{\textsf{sys}}+m)+\mathcal{C}_{\textsf{dec}}(\mathcal{L}_{\textsf{sys}}+m,\mathcal{L}_{\textsf{dec}})\Big{)}$
@@
$\lesssim n\times\frac{(\mathcal{L}_{\textsf{sys}}+m)^{2}}{m}$
$=n\times\Big{(}\frac{\mathcal{L}_{\textsf{sys}}^{2}}{m}+m+2\mathcal{L}_{\textsf{sys}}\Big{)}$
@@
The above analysis can be extended to the case with parallelism, which is especially relevant to the end-to-end latency of the overall algorithm. Considering parallel LLM calls for homogeneous sub-tasks in an ideal setting where no additional cost is induced by parallelism, we have $\mathcal{C}=\mathcal{C}(\text{sub-tasks})+\mathcal{C}(\text{aggregation})$, where
@@
$\leq\Big{(}\frac{k}{p}+1\Big{)}\times\Big{(}\mathcal{C}_{\textsf{pre}}(\mathcal{L}_{\textsf{sys}}+m)+\mathcal{C}_{\textsf{dec}}(\mathcal{L}_{\textsf{sys}}+m,\mathcal{L}_{\textsf{dec}})\Big{)}$ (5b)
@@ 3.2 Concrete examples and experiment settings
Tasks that we consider in the rest of this section include counting, sorting, retrieval, and retrieval-augmented generation (RAG), whose details will soon be explained in their corresponding subsections. For each task, we specify the concrete LLM-based algorithm, analyze its performance in terms of error and cost metrics, and validate our analysis with numerical experiments. Error metrics are task-specific, while cost metrics of interest are common among tasks, including the total prefilling length and decoding length, the total number of LLM calls, and the end-to-end latency with sequential or parallel LLM calls. These concrete examples not only confirm the practical advantages of LLM-based algorithms, but also verify that our analysis can help explain or predict the empirical performance of LLM-based algorithms, reveal the reasons behind some interesting empirical phenomena, and instruct the design of algorithms or choices of hyperparameters, e.g. the sub-task size $m$.
@@
All of these LLMs are chat models. Each LLM call involved in our algorithms is prompted in a chat format, based on the sub-task that it is responsible for. Interested readers are referred to the source code for the prompts used in our experiments.
@@ 3.5 Example: retrieval
We study a more realistic application of LLM-based algorithms, which is answering a given question that requires retrieving some key information from a long piece of text. Our design of this task draws inspiration from the needle-in-a-haystack benchmark [Kam23] and other similar benchmarks that have been widely adopted for evaluating the long-context capability of LLMs as well as techniques of retrieval-augmented generation [WBC+15, Rou23, KBA+24, ZCH+24].

### v2

@@ Abstract
We initiate a formal investigation into the design and analysis of LLM-based algorithms, i.e. algorithms that contain one or multiple calls of large language models (LLMs) as sub-routines and critically rely on the capabilities of LLMs. While LLM-based algorithms, ranging from basic LLM calls with prompt engineering to complicated LLM-powered agent systems and compound AI systems, have achieved remarkable empirical success, the design and optimization of them have mostly relied on heuristics and trial-and-errors, which is largely due to a lack of formal and analytical study for these algorithms. To fill this gap, we start by identifying the computational-graph representation of LLM-based algorithms, the design principle of task decomposition, and some key abstractions, which then facilitate our formal analysis for the accuracy and efficiency of LLM-based algorithms, despite the black-box nature of LLMs. Through extensive analytical and empirical investigation in a series of case studies, we demonstrate that the proposed framework is broadly applicable to a wide range of scenarios and diverse patterns of LLM-based algorithms, such as parallel, hierarchical and recursive task decomposition. Our proposed framework holds promise for advancing LLM-based algorithms, by revealing the reasons behind curious empirical phenomena, guiding the choices of hyperparameters, predicting the empirical performance of algorithms, and inspiring new algorithm design. To promote further study of LLM-based algorithms, we release our source code at https://github.com/modelscope/agentscope/tree/main/examples/paper_llm_based_algorithm.
@@ 1 Introduction
The rapid advancements of pre-trained large language models (LLMs) in the past few years [BCE+23, WTB+22, SMK23] have given rise to the new paradigm of algorithmic problem solving by utilizing LLMs as general-purpose solvers and prompting them with task descriptions plus additional inputs that can help enhance their performance [WWS+22, KGR+22, ZYY24]. Meanwhile, it is also well recognized that even the best LLMs today still exhibit various limitations, such as finite context window sizes and difficulty with complex reasoning. Some of these limitations are fundamental [MS23, PNP24, TTC+24, HR24, WDL24, MPS24], and resolving them requires new breakthroughs. Moreover, in resource-constrained scenarios where using the state-of-the-art LLMs is not feasible, one might have to resort to smaller and weaker LLMs for solving complex tasks.
All these have motivated the developments of what we call LLM-based algorithms, namely, algorithms that contain one or multiple LLM calls as sub-routines and fundamentally rely on the capabilities of LLMs. In its most basic form, an LLM-based algorithm can be a combination of one or multiple LLM calls and some prompt engineering [YYZ+23, BBK+23, ZSH+23, WWS+23]. More advanced examples include LLM-powered agent systems [MDL+23, PKB+24, XCG+23] and compound AI systems [ZKC+24] that augment LLMs with additional abilities like tool use and long-term memory, as well as the emerging paradigm of LLM programming [SSC+23, KSM+24, ZYX+24].
@@
  • Section 3 provides an in-depth case study for parallel decomposition, a basic pattern of task decomposition and a building block for more sophisticated LLM-based algorithms. After introducing the general formulation and analysis for this pattern, we present our algorithm design and analysis for multiple concrete tasks, which are then validated by extensive numerical experiments. Advanced case studies for hierarchical decomposition and recursive decomposition are provided in Sections 4 and 5 respectively, which further demonstrate how to apply our proposed analytical framework to diverse scenarios and complex patterns of LLM-based algorithms.
@@ 2 LLM-based algorithms
  • Other classes of problems might have limited problem sizes, but require complex reasoning for solving them correctly. In this case, LLM-based algorithms can benefit substantially from sequential task decomposition such as step-by-step thinking and acting [YZY+23], or parallel decomposition like generating multiple reasoning paths and then aggregating them for a better solution [SLC+23, WWS+23].
@@ 2.1 Algorithms as computational graphs
Given such nodes, the computational graph is built by connecting them with directed edges that specify the data flows within the LLM-based algorithm. For example, Figure 2(a) shows the graph for the pattern of parallel decomposition, which divides the input problem into parallel sub-tasks, solves each of them with one LLM node, and aggregates the results for the final solution; Figure 2(b) illustrates an algorithm for book-length summarization [CLGI24, Ope24b], which divides the input text into multiple chunks and maintains a global summary via incremental updating; and Figure 2(c) represents the ReAct algorithm [YZY+23], which consists of multiple iterations of reasoning, tool use and aggregation. The computational-graph formulation is expressive enough to cover these diverse LLM-based algorithms that have been proposed separately in prior works, and facilitates formal analysis for all of them in a unified manner. The computational graph of an LLM-based algorithm can be pre-specified and static, or dynamically constructed at runtime, as will be demonstrated later in our case studies in Sections 4 and 5. Configurations of an LLM-based algorithm include how task decomposition is done (reflected by the graph topology and the sub-tasks of graph nodes), the methods of prompting and parsing for each LLM node, configurations of the LLM(s) being used, and so on. The LLM nodes within one graph might use the same or different backbone LLMs.
@@ 2.2 Key abstractions
Error metrics are task-specific in general. Moreover, the overall algorithm and the LLM nodes within it might have different error metrics. For instance, consider an LLM-based algorithm that answers a given question by first retrieving relevant sentences from the input text with multiple LLM nodes, and then reasoning over the extracted sentences with another LLM node. In this case, the overall algorithm is evaluated by whether it answers the question correctly, while each LLM node responsible for retrieval is evaluated by whether it successfully extracts all necessary sentences within the input text that it handles.111 Throughout this work, we use “accuracy” to refer to the broader concept of “quality”, and an “error metric” can be any metric that measures how much the output of an algorithm deviates from certain criteria.
@@
Assumptions on LLM inference service [YSZ+24, PDC+23, ZNH+24], on the other hand, are task-agnostic and only affect the cost metrics, not error metrics. In particular, they determine the relation between the cost metrics and the lengths of prompt and generated text for each LLM call, the parallelism of multiple LLM calls, among others. While LLM inference service in practice can be very diverse, we will see in the next subsection that unified and formal analysis is possible with appropriate abstractions of it.
@@ 2.3 Formal analysis: accuracy and efficiency
$\mathcal{C}$ $\leq\mathcal{C}(\text{prefilling})+\mathcal{C}(\text{decoding})=\mathcal{C}_{\textsf{pre}}(\mathcal{L}_{\textsf{pre}})+\mathcal{C}_{\textsf{dec}}(\mathcal{L}_{\textsf{pre}},\mathcal{L}_{\textsf{dec}})\eqqcolon\mathcal{C}_{\textsf{LLM}}(\mathcal{L}_{\textsf{pre}},\mathcal{L}_{\textsf{dec}}).$ (2)
@@
  • Latency: in a memory-bound scenario [AQS+23] where $\mathcal{L}_{\textsf{pre}}$ is limited and the latency of memory IO for loading model weights is dominant, we have approximately $\mathcal{C}_{\textsf{pre}}(\mathcal{L}_{\textsf{pre}})=O(1)$ and $\mathcal{C}_{\textsf{dec}}(\mathcal{L}_{\textsf{pre}},\mathcal{L}_{\textsf{dec}})=O(\mathcal{L}_{\textsf{dec}})$; in more general scenarios, $\mathcal{C}_{\textsf{pre}}$ and $\mathcal{C}_{\textsf{dec}}$ might grow linearly or even quadratically (e.g. for Transformers with full attention [VSP+17]) with $\mathcal{L}_{\textsf{pre}}$ and $\mathcal{L}_{\textsf{dec}}$.
@@ 3 Case study: parallel decomposition
This section focuses on parallel decomposition, a basic MapReduce-like [DG08] pattern visualized in Figure 2(a). An algorithm of this pattern first divides the input problem into multiple independent sub-tasks, solves each of them with one LLM node, and finally aggregates the results with an LLM or non-LLM node for the final solution. The intermediate sub-tasks can be solved sequentially or in parallel, which has impacts on certain cost metrics of the overall algorithm. This basic form of task decomposition can be used as a building block for more sophisticated algorithms. Despite its simplicity, interesting analysis and a wide variety of concrete tasks and algorithms can be derived from this pattern, which are presented in the remaining of this section.
@@ 3.1 Notations and analysis
  • Finally, let $\mathcal{L}_{\textsf{sys}}$ be an upper bound for the length of the system prompt of each LLM call, which includes everything in the prompt except for the size-$n$ input problem instance. The presence of $\mathcal{L}_{\textsf{sys}}$, which can be large in practice, is essentially due to the fact that the LLM is used as a general-purpose sub-routine, hence specifying the concrete task for each LLM call constitutes part of the complexity.
@@
For notational convenience, we will often write $f(n)\lesssim g(n)$ in place of $f(n)=O(g(n))$, which means there exists a universal constant $C>0$ such that $f(n)\leq C\cdot g(n)$ for any positive integer $n$. In addition, $f(n)\asymp g(n)$ means $f(n)\lesssim g(n)$ and $g(n)\lesssim f(n)$ both hold. To further simplify notation, we will often omit the big-O notation in the input arguments of cost functions $\mathcal{C}_{\textsf{pre}}$ and $\mathcal{C}_{\textsf{dec}}$; for example, given a prompt of length $\mathcal{L}_{\textsf{sys}}+O(n)$, we will write the cost of the prefilling phase as $\mathcal{C}_{\textsf{pre}}(\mathcal{L}_{\textsf{sys}}+n)$ rather than $\mathcal{C}_{\textsf{pre}}(\mathcal{L}_{\textsf{sys}}+O(n))$.
@@
$=k\times\Big{(}\mathcal{C}_{\textsf{pre}}(\mathcal{L}_{\textsf{sys}}+m)+\mathcal{C}_{\textsf{dec}}(\mathcal{L}_{\textsf{sys}}+m,\mathcal{L}_{\textsf{dec}})\Big{)}$
@@
This basic analysis already provides some hints for tuning the hyperparameter $m$ from the perspective of minimizing costs. Below are a few possible cases, where we assume for simplicity that $\mathcal{L}_{\textsf{dec}}=O(1)$:
@@
  • It is well known that a Transformer model with full attention suffers from quadratic complexity in long-sequence processing. For this case, a reasonable assumption would be $\mathcal{C}_{\textsf{pre}}(\mathcal{L}_{\textsf{sys}}+m)+\mathcal{C}_{\textsf{dec}}(\mathcal{L}_{\textsf{sys}}+m,1)\lesssim(\mathcal{L}_{\textsf{sys}}+m)^{2}$, which implies
@@
    $\lesssim n\times\frac{(\mathcal{L}_{\textsf{sys}}+m)^{2}}{m}=n\times\Big{(}\frac{\mathcal{L}_{\textsf{sys}}^{2}}{m}+m+2\mathcal{L}_{\textsf{sys}}\Big{)}.$
@@
  • More generally, one may assume that
    $\mathcal{C}_{\textsf{pre}}(\mathcal{L}_{\textsf{sys}}+m)+\mathcal{C}_{\textsf{dec}}(\mathcal{L}_{\textsf{sys}}+m,1)\leq\alpha\times(\mathcal{L}_{\textsf{sys}}+m)^{2}+\beta\times(\mathcal{L}_{\textsf{sys}}+m)+\gamma,$
    which takes into account a more precise characterization of the FLOPs as well as memory IO for loading model weights and KV caches [AQS+23]. In this case, we have
@@
    $\leq n\times\frac{\alpha\times(\mathcal{L}_{\textsf{sys}}+m)^{2}+\beta\times(\mathcal{L}_{\textsf{sys}}+m)+\gamma}{m}$
    $=n\times\Big{(}\alpha\times m+\frac{\alpha\times\mathcal{L}_{\textsf{sys}}^{2}+\beta\times\mathcal{L}_{\textsf{sys}}+\gamma}{m}+2\times\alpha\times\mathcal{L}_{\textsf{sys}}+\beta\Big{)},$
    and the right-hand side is minimized at $m=\sqrt{\mathcal{L}_{\textsf{sys}}^{2}+\mathcal{L}_{\textsf{sys}}\times\beta/\alpha+\gamma/\alpha}$.
The above analysis can be extended to the case with parallelism, which is especially relevant to the end-to-end latency of the overall algorithm. Considering parallel LLM calls for homogeneous sub-tasks in an ideal setting where parallelism incurs no additional cost, we have $\mathcal{C}=\mathcal{C}(\text{sub-tasks})+\mathcal{C}(\text{aggregation})$, where
[…]
@@ 3.2 Concrete examples and experiment settings
Tasks that we consider in the rest of this section include counting, sorting, retrieval, retrieval-augmented generation (RAG) and long-text summarization, whose details will soon be explained in their corresponding subsections. For each task, we specify the concrete LLM-based algorithm, analyze its performance in terms of error and cost metrics, and validate our analysis with numerical experiments. Error metrics are task-specific, while cost metrics of interest are common among tasks, including the total prefilling length and decoding length, the total number of LLM calls, and the end-to-end latency with sequential or parallel LLM calls. These concrete examples not only confirm the practical advantages of LLM-based algorithms, but also verify that our analysis can help explain or predict the empirical performance of LLM-based algorithms, reveal the reasons behind some interesting empirical phenomena, and instruct the design of algorithms or choices of hyperparameters, e.g. the sub-task size $m$.
Remark 2.
While the tasks under consideration are motivated by practical scenarios, our study will mostly focus on synthetic task design, like many prior works do. This brings numerous benefits, such as avoiding data contamination, allowing full transparency and control over task configurations, and making the current work as self-contained as possible.
@@
All of these LLMs are chat models. Each LLM call involved in our algorithms is prompted in a chat format, based on the sub-task that it is responsible for. Interested readers are referred to the source code for the prompts used in our experiments. We use greedy decoding, which is deterministic, in all experiments.
@@ 3.5 Example: retrieval
We study a more realistic application of LLM-based algorithms, which is answering a given question that requires retrieving some key information from a long piece of text. Our design of this task draws inspiration from the needle-in-a-haystack benchmark [Kam23] and other similar benchmarks that have been widely adopted for evaluating the long-context capability of LLMs as well as techniques of retrieval-augmented generation [WBC+15, Rou23, KBA+24, MJ23, ZCH+24].